OpenStudy (smurfy14):

Write equation in standard form if x-intercept is 5 and y-intercept is 3.

OpenStudy (anonymous):

y=mx+b M= slope B= y intercept

OpenStudy (smurfy14):

thats not standard form though

OpenStudy (smurfy14):

standard form is Ax+By=C

OpenStudy (anonymous):

I can't remember but is the x or y in that equation considered there intercept?

OpenStudy (anonymous):

(5,0)(0,3) m=(0-3)\(5-0) m=-3/5 y-0=-3/5(x-5) y = -3x\5+3 slope intercept standard form is: 3x+5y=15

OpenStudy (anonymous):

?